# Deciphering Gas Dynamics and Star Formation in a z=1.1 Main Sequence Spiral Galaxy with ALMA and JWST

### v1

@@ 1 Introduction
With its high spatial and spectral resolution capabilities, ALMA has significantly advanced our understanding of molecular gas and dust in galaxies. By probing both the dust continuum and gas emission lines, ALMA offers detailed insights into the distribution of dust and star formation processes across diverse galactic environments and redshifts (e.g., Simpson et al., 2015; Hodge et al., 2016; Tadaki et al., 2017; Elbaz et al., 2018; Calistro Rivera et al., 2018; Hodge et al., 2019; Morishita et al., 2022; Wu et al., 2023; Lee et al., 2024b). These observations have allowed us to visualize critical components and understand their essential roles in regulating galaxy mass buildup. Additionally, the use of adaptive optics (AO) on the ground and slitless spectroscopy from space to study ionized gas emission lines has offered a comprehensive picture of galactic mass assembly (e.g., Tacchella et al., 2018; Jafariyazani et al., 2019; Wang et al., 2020). Combining these multi-wavelength observations, a physical picture emerges, explaining galaxy structural formation as a blend of extended, young star-forming disks and centrally concentrated, extremely dusty bulges.
@@
Longer wavelength IR observations enabled by JWST have made significant strides in addressing these issues by providing high-resolution images from 0.6 to 28 m. Wide-field slitless spectroscopy, now extended up to 5 m, is a powerful tool that effectively functions as a redshift machine, providing unbiased identification of strong emitters within its field of view (e.g., Sun et al., 2022, 2023; Helton et al., 2024; Morishita et al., 2024a). Additionally, its high spatial resolution opens a new window for studying spatially resolved line properties with unprecedented detail (e.g., Matharu et al., 2024; Liu et al., 2024; Naidu et al., 2024). For the first time, we are able to spatially resolve emission lines up to . traces the ionized gas and star formation activity with less sensitivity to dust attenuation due to its longer wavelength, thus serve as a robust tracer of star formation (Reddy et al., 2023). Diverse star-formation patterns revealed by WFSS-based spectra have been observed in massive galaxies, yet the physical mechanisms shaping these distinct star formation processes are not well understood (Liu et al., 2024). Understanding such diversity requires studying molecular gas and gas dynamics.
@@ 2.2 HST and JWST Grism
We utilize NIRCam WFSS F444W data from the First Reionization Epoch Spectroscopic Complete Survey (FRESCO; GO-1895; PI: P. Oesch; Oesch et al. 2023) to study emission line. The FRESCO spectra achieve a resolution of R 1600 and covers a wavelength range from 3.8 to 5.0 m, which enables us to study multiple hydrogen recombination lines across a wide range of cosmic time with a spatially resolved view. Following the methodology outlined in Sun et al. (2023)111https://github.com/fengwusun/nircam_grism/, we reduce the WFSS data using a combination of the official JWST pipeline and several customized steps. Starting with stage-1 products (rate.fits) from the MAST archive, we assign the world coordinate system (WCS) and apply flat fielding using the latest data from the JWST Calibration Reference Data System (CRDS). Background subtraction involves two steps: first, subtracting a median background created for each module and pupil, and second, using SExtractor (Bertin & Arnouts, 1996) to refine the subtraction. We extract the emission line by removing the continuum, modeled via a median filter with a smoothing kernel (Kashino et al., 2023). For a detailed description of our reduction process, we refer the reader to Liu et al. (2024).
@@ 3.1 Stellar Mass, Dust Extinction and Star Formation Rates
Dust Extinction Correction: We estimate dust extinction by using the flux ratio of Pa (from the FRESCO survey) to H (from the CLEAR survey) and comparing it with the theoretical prediction under Case B conditions. By applying the Calzetti et al. (2000) reddening curve (), we obtain the dust extinction measurement and use it to correct for extinction, thereby determining the intrinsic SFR.
@@
Star Formation Rate Calculation: The SFR is estimated by first converting the dust-corrected emission line flux to line flux, assuming the Case B recombination scenario, which predicts an intrinsic intensity ratio of / = 9.15. This approach assumes an electron temperature of K and an electron density of , as determined by photoionization models using CLOUDY version 17.02 (Ferland et al., 2017). The SFR is then estimated from the luminosity following the SFR- relation given by Kennicutt (1998) and is scaled to align with the Chabrier (2003) IMF by dividing by a factor of 1.53 (Driver et al., 2013).
For a detailed explanation of our measurements, we refer readers to Liu et al. (2024).
@@ 3.3 Dust Continuum, Gas Mass and Star Formation Efficiency
We estimate the molecular gas mass using the CO(2-1) emission line flux. As mentioned in Sec. 2.3, for the emission line flux calculation, we use only the compact array (low resolution) configuration. This decision is due to the incorrect default flux scaling of the CLEAN residual map, known as the “JvM effect” (Jorsater & van Moorsel, 1995), which may overestimate the flux in the CLEANed map when combining two configurations. The problem arises because the dirty beam resulting from the combination of two arrays is not as Gaussian as one assumes when using the clean beam. In our case, the compact array configuration already achieves a good SNR of 12.1 for calculating the total flux. Therefore, we use only this single configuration for the total flux calculation to avoid the issues associated with the JvM effect.
The velocity-integrated CO(2-1) emission line flux of our sample galaxy is Jy km s-1, consistent with the values reported in the ASPECS observations within the uncertainties (González-López et al., 2019; Aravena et al., 2019), which measured an integrated flux density of Jy km s-1 for the same target. To convert CO(2-1) to CO(1-0), we use the excitation ladder estimated by Daddi et al. (2015), where . To estimate the total molecular gas mass from the CO(1-0) luminosity, we apply a CO-to-H conversion factor (), which accounts for its dependence on gas-phase metallicity, as is known to decrease with increasing metallicity (e.g., Wilson, 1995; Leroy et al., 2011; Genzel et al., 2015; Tacconi et al., 2018). We use the metallicity-dependent calibration from Amorín et al. (2016), which gives our galaxy a metallicity of , based on the [Ne III]/[O II] ratio (see Sec. 2.2). Accordingly, we adopt for our target. As a result, the total molecular gas mass of our galaxy is . This measurement is consistent with that reported by Aravena et al. (2020) using multiple independent methods (SED-based, Rayleigh-Jeans dust-based, CO-based).
@@ 3.4 Size Measurements
We first compare the flux distribution of F090W, F444W, and the original CO(2-1) moment 0 map. The F090W image corresponds to the rest-frame 0.4 m, falling within the NUV range. The F444W image captures the rest-frame K-band ( 2 m), which is an accurate probe of the galaxy’s stellar mass (e.g., Kauffmann & Charlot, 1998; Bell & de Jong, 2001; Drory et al., 2004). Given that ASPECS-LP.3mm.06 is an extended source with a complex spiral arm structure, the intricate morphology may be adequately captured by a simple Sérsic model (Sérsic, 1963). For the NIRCam images, we deconvolve the PSF generated by WebbPSF (Perrin et al., 2014) and measure the half-light radius () from the azimuthally averaged radial profile. For the CO(2-1) map, we perform a two-dimensional elliptical Gaussian fit on the moment 0 map using the IMFIT task in CASA. The effective radii calculated for F090W, F444W, and CO are listed in Table 2.
@@ 3.5.2 Reconstructed Emission Line Map
The 2D spectra before and after reconstruction are presented in Fig. 6. We perform the same SFE calculation as described in Sec. 3.5.1, with the results displayed in Fig. 7. While the SFE obtained through the two methods shows slight differences, potentially due to uncertainties during the shift process, our conclusion remains consistent: the SFE does not change significantly across both the disk and bulge regions, regardless of the method employed.
@@ 4.1 Star Formation in Gas-Rich Disk Galaxies
The mode of star formation that feeds the growth of massive disk galaxies is still under debate. Thanks to the rich archival data set, we are able to constrain the star formation mode in ASPECS-LP.3mm.06. Its high gas fraction, moderate gas-to-star conversion efficiency, massive disk, and main-sequence level star formation activity make it an interesting sample for discussing disk galaxy growth at . In the following, we explore several scenarios that feed disk galaxy growth and their potential relation to ASPECS-LP.3mm.06.
@@ 4.2 Moderately Low Dust Extinction
Another interesting property of ASPECS-LP.3mm.06 is its modest dust reservoir. Gas-rich galaxies with active ongoing star-forming activities are generally expected to have bright FIR emission and strong dust extinction, as gas mass and dust mass are well-established to be correlated with each other at a given metallicity (Issa et al., 1990; Galametz et al., 2011; Rémy-Ruyer et al., 2014). MUSE observations using [Ne III] and [O II] reveal that ASPECS-LP.3mm.06 is a metal-rich galaxy, with . Given the high gas fraction and metallicity, it would be expected for ASPECS-LP.3mm.06 to be a dustier galaxy. However, it was undetected in the 3 mm continuum and only detected in the very deep 1.2 mm continuum (González-López et al., 2019, 2020). Furthermore, both SED-based (AV=0.8 mag) and hydrogen recombination line-based extinction values (A0.130.03 mag) suggest mild dust content. A recent study by Hodge et al. (2024) examined 13 infrared-luminous massive galaxies. One of their samples, ALESS 17.1 at , exhibits a bright dust continuum and a nearby bright spiral-like stellar structure with an offset of at the same redshift. This structure is interpreted as a merger. In contrast, our target, ASPECS-LP.3mm.06, shows no significant signs of mergers or interactions, which may be responsible for its relatively low dust content. Despite being a major population in galaxy evolution, galaxies like ASPECS-LP.3mm.06 are still underrepresented in many previous ALMA surveys due to their relatively low dust content and moderate SFRs comparable to the MS.
@@
To quantify the bulge fraction, we conduct two-component Srsic (Sérsic, 1963) fitting using galfit (Peng et al., 2002, 2010). The fitting results are presented in Fig. 10. We derive the bulge-to-total ratio from the galfit model in the F444W band, which, as discussed in Sec. 3.4, can be considered an approximate tracer of the galaxy’s stellar mass. The Srsic indices are fixed at for the bulge and for the disk components. The derived B/T ratio of 0.130.01 indicates that ASPECS-LP.3mm.06 is primarily disk-dominated but hosts a modest bulge component. This low B/T ratio suggests that the galaxy is less likely to experience morphological quenching (Martig et al., 2009; Bluck et al., 2014). Although the B/T ratio is not high enough to classify the galaxy as an early-type, the presence of a bulge remains a significant indicator of its evolutionary state. During bulge formation, the galaxy mass becomes more centrally concentrated, increasing its gravitational potential. A stronger gravitational potential forces stars and gas to move faster to remain in orbit, leading to higher velocity dispersion. This increased velocity dispersion can cause the gas to become dynamically heated or turbulent, making it difficult for the gas to collapse under its own gravity and form stars (Krumholz & McKee, 2005). This self-regulating process eventually halts star formation in the bulge as the galaxy reaches a more mature stage. ASPECS-LP.3mm.06 exhibits a significantly higher velocity dispersion ( km s-1) within the central 1 kpc, while the velocity dispersion in the disk is much lower ( km s-1), as shown in Fig. 11. As such, the elevated velocity dispersion in the bulge explains the limited efficiencies in SFE in the central region, despite the abundant gas, and also accounts for the lack of a dusty starburst that could rapidly build up the bulge. A recent study by Liu et al. (2023) utilized NIRCam and MIRI to examine the spatial distribution of PAH emission and stellar continuum in galaxies at . They found that most of their galaxies exhibited more compact PAH emission compared to the stellar continuum, suggesting active central star formation and bulge growth. In contrast, the most massive galaxy in their sample () showed similar spatial extents for both stellar and dust components, which could indicate a transition in star formation activity or the onset of quenching in the central regions. Although ASPECS-LP.3mm.06 differs from these galaxies in terms of dust content, the elevated central velocity dispersion and reduced SFE in the bulge imply that similar self-regulating processes during bulge formation may be occurring, leading to suppression of central star formation and the development of a more quiescent bulge.

### v2

@@ 1 Introduction
With its high spatial and spectral resolution capabilities, ALMA has significantly advanced our understanding of molecular gas and dust in galaxies. By probing both the dust continuum and gas emission lines, ALMA offers detailed insights into the distribution of dust and star formation processes across diverse galactic environments and redshifts (e.g., Simpson et al., 2015; Hodge et al., 2016; Tadaki et al., 2017; Elbaz et al., 2018; Calistro Rivera et al., 2018; Hodge et al., 2019; Morishita et al., 2022; Wu et al., 2023; Lee et al., 2024b). These observations have allowed us to visualize critical components and understand their essential roles in regulating galaxy mass buildup. Additionally, the use of adaptive optics (AO) on the ground (Förster Schreiber et al., 2014, 2018; Jafariyazani et al., 2019) and slitless spectroscopy from space, including observations with HST (Nelson et al., 2012, 2013, 2016; Tacchella et al., 2018; Wang et al., 2020; Simons et al., 2021; Matharu et al., 2022; Li et al., 2022) and JWST (Matharu et al., 2023; Li et al., 2023; Nelson et al., 2023; Matharu et al., 2024; Liu et al., 2024), has provided a comprehensive picture of galactic mass assembly through the study of ionized gas emission lines. Combining these multi-wavelength observations, a physical picture emerges, explaining galaxy structural formation as a blend of extended, young star-forming disks and centrally concentrated, extremely dusty bulges.
@@
Longer wavelength IR observations enabled by JWST have made significant strides in addressing these issues by providing high-resolution images from 0.6 to 28 m. Wide-field slitless spectroscopy, now extended up to 5 m, is a powerful tool that effectively functions as a redshift machine, providing unbiased identification of strong emitters within its field of view (e.g., Sun et al., 2022, 2023; Matthee et al., 2023, 2024; Helton et al., 2024; Meyer et al., 2024; Morishita et al., 2024a; Naidu et al., 2024). Additionally, its high spatial resolution opens a new window for studying spatially resolved line properties with unprecedented detail (e.g., Nelson et al., 2023; Matharu et al., 2024; Liu et al., 2024). For the first time, we are able to spatially resolve emission lines up to . traces the ionized gas and star formation activity with less sensitivity to dust attenuation due to its longer wavelength, thus serve as a robust tracer of star formation (Kennicutt & Evans, 2012; Reddy et al., 2023). Diverse star-formation patterns revealed by WFSS-based spectra have been observed in massive galaxies, yet the physical mechanisms shaping these distinct star formation processes are not well understood (Liu et al., 2024). Understanding such diversity requires studying molecular gas and gas dynamics.
@@ 2.2 HST and JWST Grism
We utilize NIRCam WFSS F444W data from the First Reionization Epoch Spectroscopic Complete Survey (FRESCO; GO-1895; PI: P. Oesch; Oesch et al. 2023) to study the emission line. The FRESCO spectra achieve a resolution of R 1600 and covers a wavelength range from 3.8 to 5.0 m, which enables us to study multiple hydrogen recombination lines across a wide range of cosmic time with a spatially resolved view. Following the methodology outlined in Sun et al. (2023)111https://github.com/fengwusun/nircam_grism/, we reduce the WFSS data using a combination of the official JWST pipeline and several customized steps. Starting with stage-1 products (rate.fits) from the MAST archive, we assign the world coordinate system (WCS) and apply flat fielding using the latest data from the JWST Calibration Reference Data System (CRDS). Background subtraction involves two steps: first, subtracting a median background created for each module and pupil, and second, using SExtractor (Bertin & Arnouts, 1996) to refine the subtraction. We extract the emission line by removing the continuum, modeled via a median filter with a smoothing kernel (Kashino et al., 2023). For a detailed description of our reduction process, we refer the reader to Liu et al. (2024).
@@ 3.1 Stellar Mass, Dust Extinction and Star Formation Rates
Dust Extinction Correction: We estimate the dust extinction at (APaα) by using the observed flux ratio of Pa (from the FRESCO survey) to H (from the CLEAR survey) and comparing it to the theoretical ratio expected under Case B recombination conditions. By applying the Calzetti et al. (2000) reddening curve (), we derive the dust extinction and use it to calculate the intrinsic flux, which is then used to determine the SFR.
@@
Star Formation Rate Calculation: The SFR is estimated by first converting the dust-corrected emission line flux to dust-corrected line flux, assuming the Case B recombination scenario, which predicts an intrinsic intensity ratio of / = 9.15. This approach assumes an electron temperature of K and an electron density of , as determined by photoionization models using CLOUDY version 17.02 (Ferland et al., 2017). The SFR is then estimated from the luminosity following the SFR- relation given by Kennicutt (1998) and is scaled to align with the Chabrier (2003) IMF by dividing by a factor of 1.53 (Driver et al., 2013).
Our calculations yield a massive galaxy with M and a star formation rate of SFR after correcting for dust extinction, with a modest dust extinction at of A mag. As shown by Reddy et al. (2023), the Calzetti et al. (2000) extinction law remains applicable even at wavelengths. Our measurement of A mag corresponds to A mag and is marginally consistent with the SED-based A reported by Aravena et al. (2020), further validating the use of the Calzetti et al. (2000) law for longer-wavelength lines such as . For a detailed explanation of our measurements, we refer readers to Liu et al. (2024).
@@ 3.3 Dust Continuum, Gas Mass and Star Formation Efficiency
We estimate the molecular gas mass (Mmol) using the CO(2-1) emission line flux. As mentioned in Sec. 2.3, for the emission line flux calculation, we use only the compact array (low resolution) configuration. This decision is due to the incorrect default flux scaling of the CLEAN residual map, known as the “JvM effect” (Jorsater & van Moorsel, 1995), which may overestimate the flux in the CLEANed map when combining two configurations. The problem arises because the dirty beam resulting from the combination of two arrays is not as Gaussian as one assumes when using the clean beam. In our case, the compact array configuration already achieves a good SNR of 12.1 for calculating the total flux. Therefore, we use only this single configuration for the total flux calculation to avoid the issues associated with the JvM effect.
The velocity-integrated CO(2-1) emission line flux of our sample galaxy is Jy km s-1, consistent with the values reported in the ASPECS observations within the uncertainties (González-López et al., 2019; Aravena et al., 2019), which measured an integrated flux density of Jy km s-1 for the same target. To convert CO(2-1) to CO(1-0), we use the excitation ladder estimated by Daddi et al. (2015), where . To estimate the total molecular gas mass from the CO(1-0) luminosity, we apply a CO-to-H conversion factor (), which accounts for its dependence on gas-phase metallicity, as is known to decrease with increasing metallicity (e.g., Wilson, 1995; Leroy et al., 2011; Genzel et al., 2015; Tacconi et al., 2018). Using the metallicity we derived for our galaxy from the [Ne III]/[O II] ratio (see Sec. 2.2), and applying the metallicity-dependent calibration from Amorín et al. (2016), we obtain . As a result, the total molecular gas mass of our galaxy is . This measurement is consistent with that reported by Aravena et al. (2020) using multiple independent methods (SED-based, Rayleigh-Jeans dust-based, CO-based).
@@ 3.4 Size Measurements
We first compare the flux distribution of F090W, F444W, and the original CO(2-1) moment 0 map. The F090W image corresponds to the rest-frame 0.4 m, falling within the NUV range. The F444W image captures the rest-frame K-band ( 2 m), which is an accurate probe of the galaxy’s stellar mass (e.g., Kauffmann & Charlot, 1998; Bell & de Jong, 2001; Drory et al., 2004).
Given that ASPECS-LP.3mm.06 is an extended source with a complex spiral arm structure that may not be adequately captured by a simple Srsic model (Sérsic, 1963), we employ a PSF deconvolution technique to measure its size in the NIRCam images, following the method of Szomoru et al. (2010), utilizing WebbPSF (Perrin et al., 2012, 2014) and Galfit (Peng et al., 2002, 2010). First, we generate the PSF for the corresponding filter using WebbPSF. We then perform Srsic fitting with Galfit to obtain the best-fit parameters by minimizing the residuals between the PSF-convolved model and the observed data. The deconvolved model is the intrinsic (PSF-free) Srsic profile generated from these best-fit parameters. By subtracting the PSF-convolved best-fit model from the observed data, we obtain residuals that capture deviations from the Srsic model. We then add these residuals back to the deconvolved Srsic model to reconstruct the intrinsic flux distribution. This combined profile is used to derive the intrinsic flux profile for size measurement. Despite residual PSF effects, this technique effectively reconstructs the true flux distribution, even for galaxies that deviate from a Srsic profile (Szomoru et al., 2010).
Finally, we measure the half-light radius () from the azimuthally averaged radial profile of the deconvolved image. To estimate the uncertainty in , we perform Monte Carlo simulations by generating 100 model galaxies. For each simulation, we randomly sample the Srsic parameters from normal distributions centered on the best-fit values obtained from Galfit, with standard deviations equal to their uncertainties. We create deconvolved Srsic models using these sampled parameters and add the corresponding residuals from the original fit to capture deviations from the Srsic model. We then measure the half-light radius for each of these modeled galaxies. The uncertainty is derived from the standard deviation of the half-light radii obtained from the simulations. By doing so, our measurement is not affected by PSF effects or the limitations of using a Srsic model to represent the complete structure.
For the CO(2–1) map, it is possible to obtain an effective radius using the same method as for JWST by deconvolving the image with the synthesized beam and modeling the deconvolved image with a two-dimensional profile. However, the image reconstruction for interferometric data depends heavily on the uv-coverage of the visibilities and the CLEANing process. These factors can introduce uncertainties in galaxy size measurements. Therefore, we measure the size directly from the visibility data using the CASA package uvmodelfit assuming Gaussian profile. Previous studies have shown that comparing image-based size measurements with visibility-based size measurements is valid and does not introduce significant systematic uncertainties (e.g., Hodge et al., 2016; Chen et al., 2020; Tadaki et al., 2020; Ikeda et al., 2022). The effective radii calculated for F090W, F444W, and CO are listed in Table 2.
@@ 3.5.2 Reconstructed Emission Line Map
The 2D spectra before and after reconstruction are presented in Fig. 6. We perform the same SFE calculation as described in Sec. 3.5.1, with the results displayed in Fig. 7. While the SFE obtained through the two methods shows slight differences, there are intrinsic distinctions between the two SFE maps. Specifically, the method using the velocity-shifted CO map and the original 2D spectra derives an SFE map that includes the galaxy’s velocity information. In contrast, the method using the original CO map and the reconstructed map derives an intrinsic SFE map without velocity information. This can lead to the star-forming clumps falling within different hexagonal bins in the resulting SFE maps, contributing to the observed differences between the two methods. Moreover, uncertainties introduced during the shifting process can also contribute to discrepancies in the SFE maps. Despite these differences, our conclusion remains consistent: the SFE does not change significantly across both the disk and bulge regions, regardless of the method employed.
@@ 4.1 Star Formation in Gas-Rich Disk Galaxies
The mode of star formation that feeds the growth of massive disk galaxies is still under debate. Thanks to the rich archival data set, we are able to constrain the star formation mode in ASPECS-LP.3mm.06. As a representative star-forming galaxy at this cosmic epoch, demonstrated by its position on the main sequence of star-forming galaxies (Fig. 9), detailed study of ASPECS-LP.3mm.06 can provide valuable insights into the physical processes governing similar galaxies at this redshift and stellar mass. In the following, we explore several scenarios that feed disk galaxy growth and their potential relation to ASPECS-LP.3mm.06.
@@ 4.2 Moderately Low Dust Extinction
Another interesting property of ASPECS-LP.3mm.06 is its modest dust reservoir. Gas-rich galaxies with active ongoing star-forming activities are generally expected to have bright FIR emission and strong dust extinction, as gas mass and dust mass are well-established to be correlated with each other at a given metallicity (Issa et al., 1990; Galametz et al., 2011; Rémy-Ruyer et al., 2014). MUSE observations using [Ne III] and [O II] reveal that ASPECS-LP.3mm.06 is a metal-rich galaxy, with . Given the high gas fraction and metallicity, it would be expected for ASPECS-LP.3mm.06 to be a dustier galaxy. However, it was undetected in the 3 mm continuum and only detected in the very deep 1.2 mm continuum (González-López et al., 2019, 2020). Furthermore, both SED-based (AV=0.8 mag) and hydrogen recombination line-based extinction values (A0.130.03 mag) suggest mild dust content. A recent study by Hodge et al. (2024) examined 13 infrared-luminous massive galaxies. One of their samples, ALESS 17.1 at , exhibits a bright dust continuum and a nearby bright spiral-like stellar structure with an offset of at the same redshift. This structure is interpreted as a merger. In contrast, our target, ASPECS-LP.3mm.06, shows no significant signs of ongoing mergers or interactions, which may be responsible for its relatively low dust content. Despite being a major population in galaxy evolution, galaxies like ASPECS-LP.3mm.06 are still underrepresented in many previous ALMA surveys due to their relatively low dust content and moderate SFRs comparable to the MS.
@@
To quantify the bulge fraction, we conduct two-component Srsic (Sérsic, 1963) fitting using galfit (Peng et al., 2002, 2010). As illustrated in Fig. 10, the central light concentration is more prominent in the F444W filter compared to F090W. In contrast, F090W reveals a greater number of small star-forming clumps, emphasizing regions of active star formation distributed across the disk. Since F444W serves as an approximate tracer of the galaxy’s stellar mass, we derive the bulge-to-total (B/T) ratio using the GALFIT model based on the F444W light distribution. In our fitting, the Srsic indices are fixed at for the bulge and for the disk components. The derived B/T ratio of 0.130.01 indicates that ASPECS-LP.3mm.06 is primarily disk-dominated but hosts a modest bulge component. This low B/T ratio suggests that the galaxy is less likely to experience morphological quenching (Martig et al., 2009; Bluck et al., 2014). Although the B/T ratio is not high enough to classify the galaxy as an early-type, the presence of a bulge remains a significant indicator of its evolutionary state. During bulge formation, the galaxy mass becomes more centrally concentrated, increasing its gravitational potential. A stronger gravitational potential forces stars and gas to move faster to remain in orbit, leading to higher velocity dispersion. This increased velocity dispersion can cause the gas to become dynamically heated or turbulent, making it difficult for the gas to collapse under its own gravity and form stars (Krumholz & McKee, 2005). This self-regulating process eventually halts star formation in the bulge as the galaxy reaches a more mature stage. ASPECS-LP.3mm.06 exhibits a significantly higher velocity dispersion ( km s-1) within the central 1 kpc, while the velocity dispersion in the disk is much lower ( km s-1), as shown in Fig. 11. As such, the elevated velocity dispersion in the bulge explains the limited efficiencies in SFE in the central region, despite the abundant gas, and also accounts for the lack of a dusty starburst that could rapidly build up the bulge. A recent study by Liu et al. (2023) utilized NIRCam and MIRI to examine the spatial distribution of PAH emission and stellar continuum in galaxies at . They found that most of their galaxies exhibited more compact PAH emission compared to the stellar continuum, suggesting active central star formation and bulge growth. In contrast, the most massive galaxy in their sample () showed similar spatial extents for both stellar and dust components, which could indicate a transition in star formation activity or the onset of quenching in the central regions. Although ASPECS-LP.3mm.06 differs from these galaxies in terms of dust content, the elevated central velocity dispersion and reduced SFE in the bulge imply that similar self-regulating processes during bulge formation may be occurring, leading to suppression of central star formation and the development of a more quiescent bulge.